# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 1.1. Acknowledgements
We thank Igor Klep, Tim Netzer, Andreas Thom, and Jurij Volčič for helpful conversations. AM is supported by NSERC Alliance Consortia Quantum grants, reference number: ALLRP 578455 - 22. WS is supported by NSERC DG 2018-03968 and an Alfred P. Sloan Research Fellowship.
@@ 2.1. Algebras and states
We work with both groups and ***-algebras. If $\mathbb{K}$ is a field with an involution, then a ***-algebra over $\mathbb{K}$ is a unital associative $\mathbb{K}$-algebra $\mathcal{A}$ with an antilinear involution $\mathcal{A}\rightarrow\mathcal{A}:a\mapsto a^{*}$ such that $(ab)^{*}=b^{*}a^{*}$ for all $a,b\in\mathcal{A}$. A ***-homomorphism $\pi:\mathcal{A}\to\mathcal{B}$ between ***-algebras $\mathcal{A}$ and $\mathcal{B}$ is an algebra homomorphism such that $\pi(a^{*})=\pi(a)^{*}$. In this paper, $\mathbb{K}$ will be either $\mathbb{C}$ or $\mathbb{Q}$ (with trivial involution), with $\mathbb{C}$ being the default choice. When $\mathcal{A}$ is a ***-algebra over $\mathbb{Q}$, we use $\mathcal{A}_{\mathbb{C}}$ to denote $\mathcal{A}\otimes_{\mathbb{Q}}\mathbb{C}$. We use $1$ for the identity in both groups and ***-algebras.
If $G$ is a group, then we let $\mathbb{K}G$ be the group algebra of $G$ over the field $\mathbb{K}$. Recall that this is a ***-algebra with $g^{*}=g^{-1}$. If $S$ is a set, then $\mathcal{F}_{S}$ will denote the free group with generating set $S$, and $\mathcal{F}_{n}:=\mathcal{F}_{\{x_{1},\ldots,x_{n}\}}$. We use the standard notation for group presentations, in which if $S$ is a set, and $R\subseteq\mathcal{F}_{S}$, then $\langle S:R\rangle$ denotes the quotient of $\mathcal{F}_{S}$ by the normal subgroup generated by $R$. We often want to add generators and relations to a given presentation $G=\langle S:R\rangle$, so if $S_{1}$ is another set disjoint from $S$, and $R_{1}\subseteq\mathcal{F}_{S\cup S_{1}}$, then we define $\langle G,S_{1}:R_{1}\rangle:=\langle S\cup S_{1}:R\cup R_{1}\rangle$. If $G$ is a group, then $G^{*n}$ denotes the free product of $n$ copies of $G$. The cyclic group of order $k$ will be denoted by $\mathbb{Z}_{k}$. If $S$ is a set, then $\mathbb{Z}_{k}^{*S}$ denotes the free product of $|S|$ copies of $\mathbb{Z}_{k}$, where the generators of the factors are denoted by the elements of $S$. In other words, $\mathbb{Z}_{k}^{*S}=\langle S:s^{k}=1\text{ for all }s\in S\rangle$.
We also define ***-algebras by means of presentations. If $S$ is a set, we let $\mathbb{K}^{*}\langle S\rangle$ denote the free noncommutative ***-algebra with generating set $S$ over a field $\mathbb{K}$. If $R\subseteq\mathbb{K}^{*}\langle S\rangle$, then $\mathbb{K}^{*}\langle S:R\rangle$ denotes the quotient of $\mathbb{K}^{*}\langle S\rangle$ by the two-sided ***-ideal generated by $R$. As with groups, if $\mathcal{A}=\mathbb{K}^{*}\langle S:R\rangle$ is a ***-algebra presentation, $S_{1}$ is a set disjoint from $S$, and $R_{1}\subseteq\mathbb{K}^{*}\langle S\cup S_{1}\rangle$, then
@@
For both group and ***-algebra presentations, we’ll follow the standard practice of writing $a=b$ for the relation $b^{-1}a$ (in the case of groups) or $a-b$ (in the case of ***-algebras). Also, a presented group (resp. algebra) is a group (resp. algebra) with a fixed presentation, and a finitely presented group (resp. algebra) is a presented group (resp. algebra) where the presentation has a finite number of generators and relations.
As mentioned above, unless specified otherwise a ***-algebra will mean a ***-algebra over $\mathbb{C}$. If $\mathcal{H}$ is a Hilbert space, then $\mathcal{B}(\mathcal{H})$ is the $C^{*}$-algebra of bounded operators on $\mathcal{H}$. A representation of a ***-algebra $\mathcal{A}$ is a ***-homomorphism $\mathcal{A}\to\mathcal{B}(\mathcal{H})$ for some Hilbert space $\mathcal{H}$. A representation is said to be finite-dimensional if $\mathcal{H}$ is finite-dimensional (we mention this notion in the preliminary but we won’t work with it in this paper). As mentioned in the introduction, a self-adjoint element $a\in\mathcal{A}$ is (representation) positive if and only if $\phi(a)$ is positive in $\mathcal{B}(\mathcal{H})$ for all Hilbert spaces $\mathcal{H}$ and representations $\phi:\mathcal{A}\to\mathcal{B}(\mathcal{H})$.
A (bounded) state on a ***-algebra $\mathcal{A}$ is a linear function $\psi:\mathcal{A}\rightarrow\mathbb{C}$ with $\psi(1)=1$ such that $\psi(a^{*})=\overline{\psi(a)}$, $\psi(a^{*}a)\geq 0$, and
@@
for all $a\in\mathcal{A}$. The GNS theorem for ***-algebras states that $\psi$ is a state if and only if there is a representation $\pi:\mathcal{A}\rightarrow\mathcal{B}(\mathcal{H})$ and a unit vector $v\in\mathcal{H}$ such that $\psi(a)=\langle v,\pi(a)v\rangle$ for all $a\in\mathcal{A}$ [Sch20, Theorem 4.38]. Given a state $\psi$, there is always such a representation $\pi$ and vector $v$ such that $\pi(\mathcal{A})v$ is dense in $\mathcal{H}$. A pair $(\pi,v)$ such that $\pi(\mathcal{A})v$ is dense in $\mathcal{H}$ and $\langle v,\pi(a)v\rangle=\psi(a)$ for all $a\in\mathcal{A}$ is called a GNS representation of $\psi$. The GNS representation theorem implies that a self-adjoint element $a\in\mathcal{A}$ is positive if and only if $f(a)\geq 0$ for all states $f$ on $\mathcal{A}$.
@@
A state $\psi$ on a ***-algebra $\mathcal{A}$ induces a semi-norm $\lVert a\rVert_{\psi}:=\sqrt{\psi(a^{*}a)}$ on $\mathcal{A}$ satisfying $\lvert\psi(ab)\rvert\leq\lVert a\rVert_{\psi}\lVert b\rVert_{\psi}$ for all $a,b\in\mathcal{A}$. A state $\psi$ on $\mathcal{A}$ is tracial if $\psi(ab)=\psi(ba)$ for all $a,b\in\mathcal{A}$. If $\psi$ is tracial, then $\lVert\cdotp\rVert_{\psi}$ is unitarily invariant, meaning that $\lVert uav\rVert_{\psi}=\lVert a\rVert_{\psi}$ for all unitaries $u,v\in\mathcal{A}$ and elements $a\in\mathcal{A}$. For an arbitrary state $\psi$, the seminorm $\lVert\cdotp\rVert_{\psi}$ is left unitarily invariant, but not necessarily right unitarily invariant.
@@
An element $a$ of a ***-algebra $\mathcal{A}$ is a hermitian square if there is $b\in\mathcal{A}$ such that $a=b^{*}b$, and a sum of (hermitian) squares if there is $b_{1},\ldots,b_{k}\in\mathcal{A}$ such that $a=\sum_{i}b_{i}^{*}b_{i}$. If $a$ is a sum of squares, then $\psi(a)\geq 0$ for all states $\psi$. If $a_{1}$ and $a_{2}$ are sums of squares, then $\psi(a_{1}a_{2})\geq 0$ for all tracial states $\psi$, since $\psi(b_{1}^{*}b_{1}b_{2}^{*}b_{2})=\psi((b_{2}b_{1}^{*})^{*}(b_{2}b_{1}^{*}))$ when $\psi$ is tracial.
Every ***-algebra $\mathcal{A}$ has an operator norm $\lVert\cdotp\rVert_{\mathcal{A}}:\mathcal{A}\rightarrow\mathbb{R}_{\geq 0}\cup\{\infty\}$ defined by $\lVert a\rVert_{\mathcal{A}}:=\sup\{\sqrt{\psi(a^{*}a)}:\psi\text{ a state on }\mathcal{A}\}$. Using the GNS theorem, it can be shown that $\lVert a\rVert_{\mathcal{A}}=\sup\{\lVert\pi(a)\rVert_{\mathcal{B}(\mathcal{H})}:\pi\text{ a }*\text{-representation}\text{ on a Hilbert space }\mathcal{H}\}$ and $\psi(b^{*}a^{*}ab)\leq\lVert a\rVert^{2}_{\mathcal{A}}\psi(b^{*}b)$ for all $a,b\in\mathcal{A}$ and states $\psi$ on $\mathcal{A}$.
@@
A ***-algebra $\mathcal{A}$ is said to be archimedean if for all $a\in\mathcal{A}$ (or equivalently, all $a$ in a generating set for $\mathcal{A}$), there is a scalar $\lambda>0$ such that $\lambda-a^{*}a$ is a sum of squares. If $\mathcal{A}$ is archimedean, then a linear functional $f:\mathcal{A}\to\mathbb{C}$ satisfying $f(1)=1$ and $f(a^{*})=\overline{f(a)}$, $f(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$ automatically satisfies the boundedness condition in Equation (2.1). Also, the operator norm $\lVert\cdotp\rVert_{\mathcal{A}}$ takes only finite values, and (as mentioned in the introduction) Helton and McCullough’s positivstellensatz states that $a\in\mathcal{A}$ is positive if and only if $a+\epsilon$ is a sum of squares for all $\epsilon>0$. Group algebras $\mathbb{C}G$ are archimedean, but free ***-algebras *$\mathbb{C}^{*}\langle S\rangle$ are not when $S$ is non-empty, so not all the algebras we deal with are archimedean. More background can be found in [Sch09] and [Oza13].
@@ 2.2. Computability
If $\Sigma$ is a set, a word of length $n$ over $\Sigma$ is a finite sequence $a_{1}a_{2}\cdots a_{n}$, where $a_{i}\in\Sigma$ for all $1\leq i\leq n$. We use the standard terminology from the theory of computation (see, e.g. [Sip12]). In particular, a subset $\mathcal{L}$ of words over $\Sigma$ is in RE if there is a Turing machine $M$ with input alphabet $\Sigma$ which recognizes $\mathcal{L}$, meaning that there is a computation path where $M$ halts on input $x$ if and only if $x\in\mathcal{L}$. In this definition $M$ can be non-deterministic, but every language in RE can be recognized by a deterministic Turing machine. A language $\mathcal{L}$ is in coRE if its complement $\overline{\mathcal{L}}$ is in RE, and is decidable if it is in both RE and coRE. Equivalently, $\mathcal{L}$ is decidable if there is a deterministic Turing machine that halts on every input $x$, returning “accept” if $x\in\mathcal{L}$ and “reject” otherwise. Since many types of objects can be encoded as strings, we often talk about sets of other objects being RE or coRE. For instance, the set
$\operatorname{HALT}=\{(M,x)\text{ a Turing machine},x\text{ an input to }M\ |\ M\text{ halts on input }x\}$
is the prototypical example of an undecidable RE set, although formalizing it as a language requires choosing some encoding of pairs $(M,x)$ as strings. The complement $\operatorname{coHALT}$ of pairs $(M,x)$ such that $M$ does not halt on input $x$ is the prototypical example of an undecidable coRE set. A function $f$ from strings on $\Sigma$ to strings on $\Sigma^{\prime}$ is computable if there is a deterministic Turing machine with input alphabet $\Sigma$ and output alphabet $\Sigma^{\prime}$ which takes input $x$ to $f(x)$. A language $\mathcal{L}$ of strings over $\Sigma$ is RE-hard (resp. coRE-hard) if there is a computable function $f$ from pairs of Turing machines $(M,x)$ to strings over $\Sigma$, such that $f(M,x)\in\mathcal{L}$ if and only if $M$ halts on input $x$ (sometimes this is called RE-hard under many-one reductions). Because natural numbers can be encoded as binary strings, we can also talk about RE, coRE, RE-hard, and coRE-hard subsets of $\mathbb{N}$.

### 18 Mar 2026 v2

@@ 1.1. Acknowledgements
We thank Igor Klep, Tim Netzer, Andreas Thom, and Jurij Volčič for helpful conversations. We thank the anonymous referee for helpful comments, including supplying reference [SS11]. AM is supported by NSERC Alliance Consortia Quantum grants, reference number: ALLRP 578455 - 22. WS is supported by NSERC DG 2018-03968 and an Alfred P. Sloan Research Fellowship.
@@ 2.1. Algebras and states
We work with both groups and *-algebras. If $\mathbb{K}$ is a field with an involution, then a *-algebra over $\mathbb{K}$ is a unital associative $\mathbb{K}$-algebra $\mathcal{A}$ with an antilinear involution $\mathcal{A}\rightarrow\mathcal{A}:a\mapsto a^{*}$ such that $(ab)^{*}=b^{*}a^{*}$ for all $a,b\in\mathcal{A}$. A *-homomorphism $\pi:\mathcal{A}\to\mathcal{B}$ between *-algebras $\mathcal{A}$ and $\mathcal{B}$ is an algebra homomorphism such that $\pi(a^{*})=\pi(a)^{*}$. In this paper, $\mathbb{K}$ will be either $\mathbb{C}$ or $\mathbb{Q}$ (with trivial involution), with $\mathbb{C}$ being the default choice. When $\mathcal{A}$ is a *-algebra over $\mathbb{Q}$, we use $\mathcal{A}_{\mathbb{C}}$ to denote $\mathcal{A}\otimes_{\mathbb{Q}}\mathbb{C}$. We use $1$ for the identity in both groups and *-algebras.
If $G$ is a group, then we let $\mathbb{K}G$ be the group algebra of $G$ over the field $\mathbb{K}$. Recall that this is a *-algebra with $g^{*}=g^{-1}$. If $S$ is a set, then $\mathcal{F}_{S}$ will denote the free group with generating set $S$, and $\mathcal{F}_{n}:=\mathcal{F}_{\{x_{1},\ldots,x_{n}\}}$. We use the standard notation for group presentations, in which if $S$ is a set, and $R\subseteq\mathcal{F}_{S}$, then $\langle S:R\rangle$ denotes the quotient of $\mathcal{F}_{S}$ by the normal subgroup generated by $R$. We often want to add generators and relations to a given presentation $G=\langle S:R\rangle$, so if $S_{1}$ is another set disjoint from $S$, and $R_{1}\subseteq\mathcal{F}_{S\cup S_{1}}$, then we define $\langle G,S_{1}:R_{1}\rangle:=\langle S\cup S_{1}:R\cup R_{1}\rangle$. If $G$ is a group, then $G^{*n}$ denotes the free product of $n$ copies of $G$. The cyclic group of order $k$ will be denoted by $\mathbb{Z}_{k}$. If $S$ is a set, then $\mathbb{Z}_{k}^{*S}$ denotes the free product of $|S|$ copies of $\mathbb{Z}_{k}$, where the generators of the factors are denoted by the elements of $S$. In other words, $\mathbb{Z}_{k}^{*S}=\langle S:s^{k}=1\text{ for all }s\in S\rangle$.
We also define *-algebras by means of presentations. If $S$ is a set, we let $\mathbb{K}^{*}\langle S\rangle$ denote the free noncommutative *-algebra with generating set $S$ over $\mathbb{K}$. Elements of $\mathbb{K}^{*}\langle S\rangle$ are noncommutative *-polynomials over $\mathbb{K}$. For instance, if $S=\{x,y\}$, then elements of $\mathbb{K}^{*}\langle S\rangle$ are noncommutative polynomials in $x,x^{*},y,y^{*}$, and the involution $a\mapsto a^{*}$ exchanges $x$ with $x^{*}$ and $y$ with $y^{*}$. If $R\subseteq\mathbb{K}^{*}\langle S\rangle$, then $\mathbb{K}^{*}\langle S:R\rangle$ denotes the quotient of $\mathbb{K}^{*}\langle S\rangle$ by the two-sided *-ideal generated by $R$. As with groups, if $\mathcal{A}=\mathbb{K}^{*}\langle S:R\rangle$ is a *-algebra presentation, $S_{1}$ is a set disjoint from $S$, and $R_{1}\subseteq\mathbb{K}^{*}\langle S\cup S_{1}\rangle$, then
@@
For both group and *-algebra presentations, we’ll follow the standard practice of writing $a=b$ for the relation $b^{-1}a$ (in the case of groups) or aba-b (in the case of *-algebras). Also, a presented group (resp. algebra) is a group (resp. algebra) with a fixed presentation, and a finitely presented group (resp. algebra) is a presented group (resp. algebra) where the presentation has a finite number of generators and relations.
As mentioned above, unless specified otherwise a *-algebra will mean a *-algebra over $\mathbb{C}$. If $\mathcal{H}$ is a Hilbert space, then $\mathcal{B}(\mathcal{H})$ is the $C^{*}$-algebra of bounded operators on $\mathcal{H}$. A representation of a *-algebra $\mathcal{A}$ is a *-homomorphism $\mathcal{A}\to\mathcal{B}(\mathcal{H})$ for some Hilbert space $\mathcal{H}$. A representation is said to be finite-dimensional if $\mathcal{H}$ is finite-dimensional (we mention this notion in the preliminary but we won’t work with it in this paper). As mentioned in the introduction, a self-adjoint element $a\in\mathcal{A}$ is (representation) positive if and only if $\phi(a)$ is positive in $\mathcal{B}(\mathcal{H})$ for all Hilbert spaces $\mathcal{H}$ and representations $\phi:\mathcal{A}\to\mathcal{B}(\mathcal{H})$.
A (bounded) state on a *-algebra $\mathcal{A}$ is a linear function $\psi:\mathcal{A}\rightarrow\mathbb{C}$ with $\psi(1)=1$ such that $\psi(a^{*})=\overline{\psi(a)}$, $\psi(a^{*}a)\geq 0$, and
@@
for all $a\in\mathcal{A}$. The GNS theorem for *-algebras states that $\psi$ is a state if and only if there is a representation $\pi:\mathcal{A}\rightarrow\mathcal{B}(\mathcal{H})$ and a unit vector $v\in\mathcal{H}$ such that $\psi(a)=\langle v,\pi(a)v\rangle$ for all $a\in\mathcal{A}$ [SCH20, Theorem 4.38]. Given a state $\psi$, there is always such a representation $\pi$ and vector $v$ such that $\pi(\mathcal{A})v$ is dense in $\mathcal{H}$. A pair $(\pi,v)$ such that $\pi(\mathcal{A})v$ is dense in $\mathcal{H}$ and $\langle v,\pi(a)v\rangle=\psi(a)$ for all $a\in\mathcal{A}$ is called a GNS representation of $\psi$. The GNS representation theorem implies that a self-adjoint element $a\in\mathcal{A}$ is positive if and only if $f(a)\geq 0$ for all states $f$ on $\mathcal{A}$.
@@
A state $\psi$ on a *-algebra $\mathcal{A}$ induces a semi-norm $\lVert a\rVert_{\psi}:=\sqrt{\psi(a^{*}a)}$ on $\mathcal{A}$ satisfying $\lvert\psi(ab)\rvert\leq\lVert a\rVert_{\psi}\lVert b\rVert_{\psi}$ for all $a,b\in\mathcal{A}$. A state $\psi$ on $\mathcal{A}$ is tracial if $\psi(ab)=\psi(ba)$ for all $a,b\in\mathcal{A}$. If $\psi$ is tracial, then $\lVert\cdotp\rVert_{\psi}$ is unitarily invariant, meaning that $\lVert uav\rVert_{\psi}=\lVert a\rVert_{\psi}$ for all unitaries $u,v\in\mathcal{A}$ and elements $a\in\mathcal{A}$. For an arbitrary state $\psi$, the seminorm $\lVert\cdotp\rVert_{\psi}$ is left unitarily invariant, but not necessarily right unitarily invariant.
@@
An element $a$ of a *-algebra $\mathcal{A}$ is a hermitian square if there is $b\in\mathcal{A}$ such that $a=b^{*}b$, and a sum of (hermitian) squares if there is $b_{1},\ldots,b_{k}\in\mathcal{A}$ such that $a=\sum_{i}b_{i}^{*}b_{i}$. If $a$ is a sum of squares, then $\psi(a)\geq 0$ for all states $\psi$. If $a_{1}$ and $a_{2}$ are sums of squares, then $\psi(a_{1}a_{2})\geq 0$ for all tracial states $\psi$, since $\psi(b_{1}^{*}b_{1}b_{2}^{*}b_{2})=\psi((b_{2}b_{1}^{*})^{*}(b_{2}b_{1}^{*}))$ when $\psi$ is tracial.
[…]
@@
A *-algebra $\mathcal{A}$ is said to be archimedean if for all $a\in\mathcal{A}$ (or equivalently, all $a$ in a generating set for $\mathcal{A}$), there is a scalar $\lambda>0$ such that $\lambda-a^{*}a$ is a sum of squares. If $\mathcal{A}$ is archimedean, then a linear functional $f:\mathcal{A}\to\mathbb{C}$ satisfying $f(1)=1$ and $f(a^{*})=\overline{f(a)}$, $f(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$ automatically satisfies the boundedness condition in Equation (2.1). Also, the operator norm $\lVert\cdotp\rVert_{\mathcal{A}}$ takes only finite values, and (as mentioned in the introduction) Helton and McCullough’s Positivstellensatz states that $a\in\mathcal{A}$ is positive if and only if $a+\epsilon$ is a sum of squares for all $\epsilon>0$. Group algebras $\mathbb{C}G$ are archimedean, but free *-algebras $\mathbb{C}^{*}\langle S\rangle$ are not when $S$ is non-empty, so not all the algebras we deal with are archimedean. More background can be found in [SCH09] and [OZA13].
@@ 2.2. Computability
If $\Sigma$ is a set, a word (also called a string) of length $n$ over $\Sigma$ is a finite sequence $a_{1}a_{2}\cdots a_{n}$, where $a_{i}\in\Sigma$ for all $1\leq i\leq n$. We use the standard terminology from the theory of computation (see, e.g. [SIP12]). In particular, a subset $\mathcal{L}$ of words over a finite set $\Sigma$ is in RE if there is a Turing machine $M$ with input alphabet $\Sigma$ which recognizes $\mathcal{L}$, meaning that there is a computation path where $M$ halts on input $x$ if and only if $x\in\mathcal{L}$. In this definition $M$ can be non-deterministic, but every language in RE can be recognized by a deterministic Turing machine. A language $\mathcal{L}$ is in coRE if its complement $\overline{\mathcal{L}}$ is in RE, and is decidable if it is in both RE and coRE. Equivalently, $\mathcal{L}$ is decidable if there is a deterministic Turing machine that halts on every input $x$, returning “accept” if $x\in\mathcal{L}$ and “reject” otherwise. Since many types of objects can be encoded as strings, we often talk about sets of other objects being RE or coRE. For instance, the set
[…]
is the prototypical example of an undecidable RE set, although formalizing it as a language requires choosing some encoding of pairs $(M,x)$ as strings. The complement $\operatorname{coHALT}$ of pairs $(M,x)$ such that $M$ does not halt on input $x$ is the prototypical example of an undecidable coRE set. A function $f$ from strings on $\Sigma$ to strings on $\Sigma^{\prime}$ is computable if there is a deterministic Turing machine with input alphabet $\Sigma$ and output alphabet $\Sigma^{\prime}$ which takes input $x$ to $f(x)$. A language $\mathcal{L}$ of strings over $\Sigma$ is RE-hard (resp. coRE-hard) if there is a computable function $f$ from pairs of Turing machines $(M,x)$ to strings over $\Sigma$, such that $f(M,x)\in\mathcal{L}$ if and only if $M$ halts on input $x$ (sometimes this is called RE-hard under many-one reductions). The definition of RE and coRE does not depend on the choice of alphabet $\Sigma$, because strings over a finite set can be reencoded as strings over $\{0,1\}$. Strings over $\{0,1\}$ are called bit strings or binary strings. Natural numbers can be encoded as binary strings as well, so we can also talk about RE, coRE, RE-hard, and coRE-hard subsets of $\mathbb{N}$.
@@
The classes RE and coRE are the first level of a larger hierarchy of classes called the arithmetical hierarchy. Although we don’t use it in the rest of the paper, the class $\Pi_{2}^{0}$ mentioned in the introduction is in the second level of this hierarchy. A language $\mathcal{L}$ is in $\Pi_{2}^{0}$ if there is a Turing machine $M$ taking a pair of strings $x$ and $y$ as input, such that $x\in\mathcal{L}$ if and only if $M$ halts on input $(x,y)$ for all $y$. As discussed in the introduction, the Helton-McCullough Positivstellensatz states that an element $\alpha$ of an archimedean *-algebra $\mathcal{A}$ is positive if and only if $\alpha+\epsilon$ is a sum of squares for all $\epsilon>0$. When $\mathcal{A}$ is finitely presented, one can effectively search through all sums of squares. So deciding positivity for finitely presented archimedean algebras is in $\Pi_{2}^{0}$.